[Brahms-l] Shift report 20050429 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Apr 29 2005 - 07:56:38 EDT
Shifter leader was SB
Shifters were 


1. First store had to be dumped during machine setup since C-AD could not maintain the beam. Second store since 5:46 am is stable and is scheduled to be dumped at 11:45 am.

2. Fs@8 deg @2442A,  MRS@45 deg @1050A
trig2(3.7k), trig3(314k), trig4(56k),trig5(432k),trig6(118k)

3. No detector problems observed

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Apr 29 07:57:11 2005

This archive was generated by hypermail 2.1.8 : Fri Apr 29 2005 - 07:57:23 EDT